what the answer for this question (3)+(2)+(-1)
Answer by Alan3354(69443)

what the answer for this question (3)+(2)+(-1)
---------
There's no question.
If you want to add those:
(3)+(2)+(-1) = 3+2-1
= 5-1
= 4